# Adversarial-MidiBERT: Symbolic Music Understanding Model Based on Unbias Pre-training and Mask Fine-tuning

### v1

@@ Abstract
As an important part of Music Information Retrieval (MIR), Symbolic Music Understanding (SMU) has gained substantial attention, as it can assist musicians and amateurs in learning and creating music. Recently, pre-trained language models have been widely adopted in SMU because the symbolic music shares a huge similarity with natural language, and the pre-trained manner also helps make full use of limited music data. However, the issue of bias, such as sexism, ageism, and racism, has been observed in pre-trained language models, which is attributed to the imbalanced distribution of training data. It also has a significant influence on the performance of downstream tasks, which also happens in SMU. To address this challenge, we propose Adversarial-MidiBERT, a symbolic music understanding model based on Bidirectional Encoder Representations from Transformers (BERT). We introduce an unbiased pre-training method based on adversarial learning to minimize the participation of tokens that lead to biases during training. Furthermore, we propose a mask fine-tuning method to narrow the data gap between pre-training and fine-tuning, which can help the model converge faster and perform better. We evaluate our method on four music understanding tasks, and our approach demonstrates excellent performance in all of them. The code for our model is publicly available at https://github.com/RS2002/Adversarial-MidiBERT.
@@ 1 Introduction
Music Information Retrieval (MIR) plays a crucial role in various fields, such as the recommendation systems in music apps and the AI agents for music creation. With the advancement of computer music, symbolic music, which represents music through a structural sequence of notes, has gained widespread attention because most current music is initially created and recorded using symbolic music formats like MIDI [1]. Symbolic Music Understanding (SMU) has been a key research direction within MIR, aiming to assist musicians and amateurs in learning, teaching, and creating music.
@@
Given the similarity between symbolic music and natural language, language models have been widely used in SMU. For example, the Bidirectional Encoder Representations from Transformers (BERT) [2] model has shown promising performance in SMU [3, 4]. An important factor in the success of current language models, especially Large Language Models (LLMs), is their use of large amounts of unlabeled data to pre-train the models to learn basic data structure and relationships. This pre-training mechanism has also been effective in domains with limited data, such as music [5] and signals [6], which can help improve model performance in downstream tasks [7].
Currently, the most popular pre-training method in language model is Mask Language Model (MLM) [2], which is also widely used in pre-trained SMU models [3, 4, 5]. However, due to dataset imbalances, it can lead to bias problems like discrimination in sexism, ageism, and racism in the field of Natural Language Processing (NLP) [8, 9]. For example, in the sentence “She is good at math.”, the MLM model would randomly mask some tokens and train the model to recover them. If the masked sentence becomes “[MASK] is good at math.”, the model can only recover it according to the training data distribution to achieve the highest accuracy, instead of considering the contextual relationship, because there is no information pointing to the subject. When the training set data distribution about gender in different context scenarios is imbalanced, the model may prefer to recover the [MASK] as “He”, resulting in a gender bias problem.
Recently, some studies have indicated that the bias problem can also significantly harm the model performance in downstream tasks including classification and generation [9]. The MLM-based pre-trained models in other areas also suffer from the same problem since the context-free tokens can only be recovered by the data distribution. However, most current methods to solve this problem are limited to the field of NLP and are difficult to transfer to other areas. Most of these methods can only solve single bias problems like sexism [10] and region [11], but other areas do not have a similar concept. For example, data augmentation [12] is a promising method in NLP, but due to the bias problem in other fields like music and signal not being as clear as in natural language, we do not know how to effectively clean, generate or modify the data without sufficient domain-specific knowledge.
To address the bias problem in SMU, we first need to consider what we want the model to learn during MLM pre-training. It should be the music structure, relationships, and regularities like basic mode, riff regularities, and modulation regularities, rather than the harmony, chord, or melody direction within specific styles, where those domain-based knowledge would also influence the model performance when the test set has a significant domain gap with the training set. In other words, we want the model to acquire context-dependent information rather than relying solely on the training set’s data distribution.
@@
Based on this premise, we propose Adversarial-MidiBERT for SMU: (1) We design an adversarial mechanism to try avoiding masking those context-free tokens to mitigate the bias problem. (2) We propose a mask fine-tuning method that adds random [MASK] tokens during fine-tuning to narrow the data gap between pre-training and fine-tuning. This approach efficiently improves the convergence speed and model performance. (3) Experimental results show that our method achieves excellent performance on four music understanding tasks, including composer classification, emotion classification, velocity prediction, and melody extraction.
@@ 2.1 Overview
The main structure of our Adversarial-MidiBERT is illustrated in Fig. 1. It takes BERT [2] (the encoder part of the Transformer [13]) as the backbone, whose bi-directional attention mechanism can efficiently capture the relationships in music. To adapt BERT for the SMU task, we modify the bottom embedding layer to encode music information and the top output heads for pre-training and downstream tasks.
To embed MIDI music information, we first employ Octuple [3] to represent the symbolic music structure. It transfers each MIDI file into a sequence of tokens, where each token has eight attributes: time signature (TS), tempo (BPM), bar position (BAR), relative position within each bar (POS), instrument, pitch, duration, and velocity. We then use eight embedding layers to encode these eight attributes respectively and concatenate them together.
As for the top layer of our model, there are four different heads: masker, recoverer, token-level classifier, and sequence-level classifier, all of which share the same backbone. During pre-training, we use the masker to generate the probability of masking each token, and the recoverer would recover the masked tokens. During fine-tuning, we use the classifier heads for classification tasks. The sequence-level classifier generates a single label for a whole music piece, useful for tasks like composer classification and emotion classification. The token-level classifier generates a label sequence corresponding to each token, useful for tasks like velocity prediction and melody extraction.
To address the bias problem in pre-training, we design an adversarial learning mechanism. The masker tries to mask tokens that are difficult for the recoverer to recover, while the recoverer tries to recover all masked tokens. After several epochs, the masker selects context-free tokens with the highest probability, as they can only be inferred according to the training data distribution, leading to the lowest recovery accuracy. We then freeze these tokens to prevent the recoverer from masking them in subsequent epochs. We also apply an unfreezing mechanism, randomly unfreezing some frozen tokens to avoid incorrect freezing. This adversarial process continues until the model converges.
During fine-tuning, we design a mask fine-tuning mechanism, where random [MASK] tokens replace input tokens to reduce the gap between pre-training and fine-tuning. This approach improves convergence speed and model performance.
2.2 Unbias Pre-train
@@
$L_{masker}=\sum_{i\in I_{0}}\text{MSE}(p_{i},0)+\sum_{i\in I_{1}}\text{MSE}(p_{i},1)\ ,$ (3)
@@
2.3 MASK Fine-tune
@@ 3.2 Experiment Result
Table 3: Model Performance in Different Tasks: The bold and underlined value indicates the best and second best result within each task.
Model Pre-train Sequence-Level Classification Token-Level Classification Accuracy Epochs Time Composer Emotion Velocity Melody MidiBERT [4] 79.60% 500 6.44d 79.07% 67.59% 44.88% 92.53% MusicBERT-QM [20] 80.57% 500 9.47d 83.72% 69.52% 46.71% 92.64% MusicBERT [3] 76.01% 500 10.06d 86.05% 71.06% 38.79% 92.47% PianoBART [5] 96.67% 268 3.19d 88.37% 73.15% 49.37% 92.62% Adversarial-MidiBERT (ours) 81.47% 436 9.82d 97.92% 79.46% 45.58% 92.68% Adversarial-MidiBERT (fine-tune w/o mask) 81.47% 436 9.82d 65.98% 70.53% 45.30% 92.55% Adversarial-MidiBERT (w/o pre-train) 79.76% 68.75% 38.70% 87.98%
In this section, we compare our method with other SMU models, including MidiBERT [4], MusicBERT [3], MusicBERT-QM [20], and PianoBART [5]. The main differences between these methods lie in their pre-training approaches. For fairness, these BERT-based models use the same backbone as our method. The experimental results are shown in Table 3. It can be seen that our method outperforms the previous BERT-based methods in most tasks, but loses to PianoBART in pre-training and velocity prediction. This may be influenced by the model structure, as the auto-regressive mechanism of the decoder structure in Bidirectional and Auto-Regressive Transformers (BART) [21] makes it more suitable for sequence-level tasks. However, it can be noticed that our model has an extremely significant increase in performance on sequence-level tasks.
We also conducted an ablation study to illustrate our method’s performance without pre-training and without masking during fine-tuning. The results show that the model’s performance decreases to varying degrees in these cases, demonstrating the effectiveness of our proposed mechanisms. Additionally, we observe that our model has a faster convergence speed than others. By the end of the first epoch, our method achieves relatively high accuracy in all downstream tasks.
@@ 4 Conclusion
In this paper, we present Adversarial-MidiBERT for SMU, the first method to address the bias problem of pre-trained models in MIR through adversarial pre-training method. Additionally, we introduce a mask fine-tuning approach that significantly enhances the model’s accuracy and convergence speed on downstream tasks. Our method achieves remarkable performance on four SMU tasks, especially on the sequence-level tasks. In the future, we aim to explore the application of our method to music generation tasks and NLP tasks.

### v2

# Let Network Decide What to Learn: Symbolic Music Understanding Model Based on Large-scale Adversarial Pre-training

@@ Abstract
As a crucial aspect of Music Information Retrieval (MIR), Symbolic Music Understanding (SMU) has garnered significant attention for its potential to assist both musicians and enthusiasts in learning and creating music. Recently, pre-trained language models have been widely adopted in SMU due to the substantial similarities between symbolic music and natural language, as well as the ability of these models to leverage limited music data effectively. However, some studies have shown the common pre-trained methods like Mask Language Model (MLM) may introduce bias issues like racism discrimination in Natural Language Process (NLP) and affects the performance of downstream tasks, which also happens in SMU. This bias often arises when masked tokens cannot be inferred from their context, forcing the model to overfit the training set instead of generalizing. To address this challenge, we propose Adversarial-MidiBERT for SMU, which adaptively determines what to mask during MLM via a masker network, rather than employing random masking. By avoiding the masking of tokens that are difficult to infer from context, our model is better equipped to capture contextual structures and relationships, rather than merely conforming to the training data distribution. We evaluate our method across four SMU tasks, and our approach demonstrates excellent performance in all cases. The code for our model is publicly available at https://github.com/RS2002/Adversarial-MidiBERT.
@@ 1 Introduction
Music Information Retrieval (MIR) plays a crucial role in various applications, including recommendation systems in music apps and AI agents for music creation. With advancements in computer music, symbolic music—representing music through a structured sequence of notes—has gained significant attention, as most contemporary music is initially created and recorded using symbolic formats like MIDI [1]. Symbolic Music Understanding (SMU) has emerged as a key research direction within MIR, aiming to assist both musicians and enthusiasts in learning, teaching, and creating music.
@@
Given the similarities between symbolic music and natural language, language models have been extensively utilized in SMU. For instance, the Bidirectional Encoder Representations from Transformers (BERT) [2] model has demonstrated promising performance in SMU [3, 4]. A critical factor contributing to the success of current language models, particularly Large Language Models (LLMs), is their ability to leverage vast amounts of unlabeled data during pre-training to learn fundamental data structures and relationships. This pre-training mechanism has proven effective in domains with limited data, such as music [5] and signals [6], thereby enhancing model performance in downstream tasks [7].
Among the prevalent pre-training methods for language models is the Mask Language Model (MLM) [2], particularly in encoder-only architectures. This method randomly masks certain tokens and trains the model to recover them. However, since some tokens cannot be inferred from the given context, training the model on such tasks can introduce bias issues, such as discrimination based on gender, age, or race in Natural Language Processing (NLP) [8]. Research has indicated that these biases can significantly impair model performance in downstream tasks, including classification and generation [9]. Consequently, even in the absence of ethical discrimination issues in music, bias can adversely affect model performance in SMU.
Before delving deeper into the bias problem, we present a straightforward example from NLP that may be more accessible to most readers than a music-related example. Consider the original sentence, ”She is good at math.” If we mask it to read, ”[MASK] is good at math.”, the model can only recover the masked token based on the training data distribution for optimal accuracy, rather than considering the contextual relationship, as there is no information indicating the subject. When the training data distribution regarding gender in various contexts is imbalanced, the model may default to recovering the [MASK] as ”He”, resulting in gender bias. In this case, we hope the model learns the basic grammatical structure of phrases like ”be good at”, rather than inferring ”she” and ”math” solely based on data distribution to achieve higher accuracy. Similarly, in music, we aim for the model to learn common musical structures, relationships, and regularities—such as basic modes, riff patterns, and modulation techniques—rather than developing habits specific to particular composers. In this paper, we define tokens that can be inferred from context as context-dependent tokens and those that cannot as context-free tokens. We argue that only masking context-dependent tokens in MLM is beneficial, as training the model to recover context-free tokens may lead to overfitting to the training set.
However, most current methods addressing bias are limited to the field of NLP and are challenging to transfer to other domains. Many of these methods tackle specific bias issues, such as sexism [10] and regional bias [11], but do not apply similarly in other areas. For instance, data augmentation [12] is a promising approach in NLP. However, due to the less pronounced bias issues in domains like music and signal processing, it remains unclear how to effectively clean, generate, or modify data without sufficient domain-specific knowledge.
@@
To address the bias problem in SMU, we propose the Adversarial-MidiBERT: (1) We design an adversarial mechanism to avoid masking context-free tokens in order to mitigate bias. (2) We introduce a mask fine-tuning method that incorporates random [MASK] tokens during fine-tuning to bridge the data gap between pre-training and fine-tuning, thereby enhancing model performance. (3) Experimental results demonstrate that our method achieves excellent performance across four music understanding tasks, including composer classification, emotion classification, velocity prediction, and melody extraction.
@@ 2.1 Overview
Similar to most BERT-based methods, our Adversarial-MidiBERT involves two phases: MLM pre-training and fine-tuning. During pre-training, we introduce an adversarial mechanism to prevent context-free tokens from being masked. Specifically, we train a masker network to select tokens for masking and a recoverer network to recover them. The masker aims to minimize the recovery accuracy of the recoverer, thereby favoring the selection of context-free tokens. These tokens can only be inferred from the dataset distribution, which does not guarantee precise recovery. It is likely that the recovery accuracy for context-free tokens will be lower than that for context-dependent tokens, as a well-trained recoverer can easily infer context-dependent tokens from the surrounding information. After several epochs, we can freeze the tokens with a high selection probability according to the masker, preventing them from being masked in subsequent training. This allows the recoverer to focus solely on learning from context-dependent tokens. During fine-tuning, we implement a mask fine-tuning mechanism, where random [MASK] tokens replace input tokens to reduce the gap between pre-training and fine-tuning, thereby improving model performance efficiently.
Before delving deeper into the training mechanism, we first introduce our network and data structure. To embed MIDI music information, we employ Octuple [3] to represent the symbolic music structure. It transforms each MIDI file into a sequence of tokens, each possessing eight attributes: time signature (TS), tempo (BPM), bar position (BAR), relative position within each bar (POS), instrument, pitch, duration, and velocity. To input these tokens into the network, we utilize eight embedding layers to embed them individually, and then concatenate the embedding results. This process yields an embedding sequence suitable for input to the network, as shown in Fig. 1. In this configuration, a BERT [2] (the encoder component of the Transformer [13]) serves as the backbone, complemented by an embedding layer for music information and multiple output heads for different tasks. Specifically, the recoverer head addresses the MLM task by generating classification probabilities for each attribute at each position. The masker head generates the mask probability for each token. Additionally, the sequence-level and token-level classifiers are employed in downstream classification tasks, producing classification probabilities for the entire music sequence and each token, respectively. Due to page limitations, please refer to [5, 3] for more details regarding the model structure and Octuple information.
@@ 2.2 Adversarial Pre-training
$L_{masker}=\sum_{i\in I_{0}}\text{BCE}(p_{i},0)+\sum_{i\in I_{1}}\text{BCE}(p_{i},1)\ ,$ (3)
@@
2.3 MASK Fine-tuning
@@ 3.1 Experiment Setup
Table 3: Model Performance in Different Tasks: The bold and underlined value indicates the best and second best result within each task. (Note that the training times provided are for reference only, as the server is shared with other users.)
Model Pre-train Sequence-Level Classification Token-Level Classification Accuracy Epochs Time Composer Emotion Velocity Melody MidiBERT [4] 79.60% 500 6.44d 79.07% 67.59% 44.88% 92.53% MusicBERT-QM [20] 80.57% 500 9.47d 83.72% 69.52% 46.71% 92.64% MusicBERT [3] 76.01% 500 10.06d 86.05% 71.06% 38.79% 92.47% PianoBART [5] 96.67% 268 3.19d 88.37% 73.15% 49.37% 92.62% Adversarial-MidiBERT (ours) 81.47% 436 9.82d 97.92% 79.46% 45.58% 92.68% Adversarial-MidiBERT (fine-tune w/o mask) 81.47% 436 9.82d 65.98% 70.53% 45.30% 92.55% Adversarial-MidiBERT (pre-train w/o adversary) 83.51% 500 5.93d 88.91% 74.07% 45.44% 92.64% Adversarial-MidiBERT (w/o pre-train) 79.76% 68.75% 38.70% 87.98%
@@ 3.2 Experiment Result
In this section, we compare our method with other SMU models, including MidiBERT [4], MusicBERT [3], MusicBERT-QM [20], and PianoBART [5]. The main differences between these methods lie in their pre-training approaches. For fairness, these BERT-based models use the same backbone as our method. The experimental results are shown in Table 3. It can be seen that our method outperforms the previous BERT-based methods in most tasks, but loses to PianoBART in pre-training and velocity prediction. This may be influenced by the model structure, as the auto-regressive mechanism of the decoder structure in Bidirectional and Auto-Regressive Transformers (BART) [21] makes it more suitable for token-level tasks. However, it can be noticed that our model has an extremely significant increase in performance on sequence-level tasks. What’s more, we also conducted an series of ablation study to illustrate our method’s performance without pre-training, pre-training without the adversarial-learning mechanism and without masking during fine-tuning. The results show that the model’s performance decreases to varying degrees in these cases, demonstrating the effectiveness of our proposed mechanisms.
@@ 4 Conclusion
In this paper, we present Adversarial-MidiBERT for SMU, addressing the bias problem of pre-trained models in MIR through an adversarial pre-training method. Additionally, we introduce a mask fine-tuning approach that significantly enhances the model’s accuracy on downstream tasks. Our method achieves remarkable performance on four SMU tasks, particularly on sequence-level tasks.